OpenStudy (anonymous):

find the derivative of 8y=3x^5-5Ox^3+135x

OpenStudy (lgbasallote):

use power rule \[\frac d{dx} (8y) \implies 8y^{\prime}\] \[\frac d{dx} (3x^5) \implies 5 \times 3x^{5 - 1} \implies 15x^4\] \[\frac d {dx} (50x^3) \implies 3 \times 50x^{3-1} \implies 150x^2\] \[\frac d{dx} (135x) \implies 135\] does that help?

OpenStudy (anonymous):

i will use derivative of a quotient for that solution or not? please answer me.

OpenStudy (lgbasallote):

quotient rule? no. this is purely power rule

OpenStudy (lgbasallote):

although...you'll have to divide both sides by the coefficient of y' later on to isolate y'

OpenStudy (anonymous):

thanks a lot. i'll try to solve it now. ^____^

OpenStudy (anonymous):

@Igbasallote because i was ask to get the roots i got \[8y \prime=15(x^2-9)(x^2-1)\] will become \[8y \prime= 15 (x-3)(x+3)(x-1)(x+1)\] how can i get the roots if there is 8 in y'?

OpenStudy (lgbasallote):

divide both sides by 8

OpenStudy (anonymous):

only that step? so it will be disregarded if ever and the roots are 3, -3, 1, and -1.is that so?

OpenStudy (lgbasallote):

pretty much. constant factors are always disregarded

OpenStudy (anonymous):

thanks i get now

OpenStudy (lgbasallote):

wonderful